• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 00:04:34 on June 05, 2005

    Entry number 145057

    keyword=Swing shift summary

    Run Crew: Brad Sawatzky, Ramesh Subedi

    Notable Events:

    ---------------

    1600 Problem with coolant flow to hall caused most of the L/RHRS magnets to trip at 1550. Recovery in progress.

    1700 LHRS and septum magnets OK, RHRS still ramping down

    1710 Started taking some optics data using LHRS (see run summary)

    1915-2030 Completed the magnet cross-talk test portion of the optics runplan (runs 4238--4244)

    2025 RHRS back to set-points (1.850 GeV)

    2040 Started Momentum scan portion of optics run plan (run 4245--)

    Run Summary:

    --------------------------

    4428 Thin C, RHRS ramping down, LHRS good, raster OFF, 500k

    4429 Thin C, RHRS ramping down, LHRS good, raster OFF, 500k

    4430 Thin C, RHRS ramping down, LHRS good, raster OFF, 500k

    4431 Thick C, RHRS ramping down, LHRS good, raster OFF, 500k

    4432 Waterfall, RHRS ramping up, LHRS good, raster OFF, 500k

    4433 Waterfall, RHRS ramping up, LHRS good, raster OFF, 1M

    4434 Waterfall, RHRS ramping up, LHRS good, raster OFF, 1M

    -- (Raster at 2x2)

    4435 Target at Waterfall position but water OFF, raster ON, RHRS ramping up, LHRS good 1M

    4436 Target at Waterfall position but water ON, raster ON, RHRS ramping up, LHRS good 1M

    -- Magnet cross-talk runs

    4437 Thick C, raster OFF, RHRS ramping up, LHRS good, R Septum ON

    4438 Thick C, raster OFF, RHRS ramping up, LHRS good, R Septum OFF

    4439 Thick C, raster OFF, RHRS ramping up, LHRS good, R Septum OFF, RQ1 OFF

    4441 Thick C, raster OFF, RHRS ramping up, LHRS good, R Septum OFF, RQ1 positive 50%

    4442 Thick C, raster OFF, RHRS ramping up, LHRS good, R Septum OFF, RQ1 positive 100%

    4443 Thick C, raster OFF, RHRS good, LHRS good, R Septum ON, RQ1 negative polarity again, All magnets set for 1.850 GeV/c

    4444 Thick C, raster OFF, All magnets set for 1.850 GeV/c

    -- Optic Momentum Scan

    4445 Thin C, raster OFF, PO=1.850

    4446 Ta, raster OFF, PO=1.850

    4447 Waterfall, raster OFF, PO=1.850

    4448 Waterfall, raster ON, PO=1.850

    4449 Thin C, raster ON, PO=1.850

    4450 Thin C, raster OFF, PO=1.924 --- JUNK septum not changed

    4451 Ta, raster OFF, PO=1.924 --- JUNK septum not changed

    4452 Ta, raster OFF, PO=1.887 --- JUNK septum not changed

    4453 Thin C, raster OFF, PO=1.887 --- JUNK septum not changed